# Jarník-Besicovitch type theorems for semisimple algebraic groups

### v1

@@ 4. A lower bound for the Hausdorff dimension of $S_{\gamma}^{c}$
$a_{-t_{k}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}}\subset B_{\mathbf{F}_{e}}(\tilde{\delta}_{0},T_{k}).$
@@
if $T_{k}>0$ and $l_{k}>0$ are sufficiently large. Note that $\Phi^{0}(\mathbf{F}_{e})$ may not be empty, and some of these cubes may be outside $S$, and here we collect only those cubes which are inside $S$. Thus, we obtain a family of disjoint cubes constructed from sets in $\mathcal{F}_{k}(S)$ inside $S$. We denote by $\mathcal{A}_{k}$ the collection of all these disjoint cubes constructed from $\mathcal{F}_{k}(S)$ inside $S$ where $S$ ranges over all elements in $\mathcal{A}_{k-1}$.
In this manner, we obtain a sequence $\left\{l_{k}\right\}_{k\in\mathbb{N}}$ of sufficiently large numbers $l_{k}$ with
$l_{k+1}\gg l_{k}\;(\forall k\in\mathbb{N})$
and a tree-like structure $\left\{\mathcal{A}_{k}\right\}_{k\in\mathbb{N}}$ of finite collections of cubes. Using the notation in Theorem 4.3, we have
$d_{k}(\mathcal{A})\sim l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
@@
$\begin{cases}\Delta_{k}(\mathcal{A})\sim l_{k+1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\beta_{0}(a_{1})}}\cdot\prod_{\alpha\in\Phi(\mathbf{F}_{e})}l_{k+1}^{\frac{\alpha(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}},&\Phi^{0}(\mathbf{F}_{e})=\emptyset\\ \Delta_{k}(\mathcal{A})\sim l_{k+1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\beta_{0}(a_{1})}}\cdot\prod_{\alpha\in\Phi^{1}(\mathbf{F}_{e})}l_{k+1}^{\frac{\alpha(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}\cdot\prod_{\alpha\in\Phi^{0}(\mathbf{F}_{e})}l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}},&\Phi^{0}(\mathbf{F}_{e})\neq\emptyset.\end{cases}$
Now let $\mathbf{A}_{\infty}=\bigcap_{k\in\mathbb{N}}\mathbf{A}_{k}$. By applying Theorem 4.3, we can compute (assuming that $l_{k+1}$ is much larger than $l_{k}$ for any $k\in\mathbb{N}$) that
@@
$=$ $\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}.$
@@
We have $\mathbf{A}_{\infty}\subset S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
@@
By the construction, for any $p\in\mathbf{A}_{\infty}$, there exists a sequence of rational points
@@
Note that by definition, the denominator of the rational element $a_{t_{k}}\cdot q_{k}$ is equal to the co-volume of the lattice $a_{t_{k}}\cdot q_{k}\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}$ and
$d(a_{t_{k}}\cdot q_{k})=e^{\beta_{0}(a_{t_{k}})\cdot\dim V_{\beta_{0}}}\cdot d(q_{k}).$
By Lemma 2.6, we deduce that
$\delta(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})$ $\sim\delta(\rho(a_{t_{k}}\cdot q_{k})\mathbb{Z}^{d})\lesssim d(a_{t_{k}}\cdot q_{k})^{\frac{1}{\dim V_{\beta_{0}}}}$
$\lesssim e^{\beta_{0}(a_{t_{k}})}\cdot l_{k}\leq e^{-(\tau+\epsilon)t_{k}}$
and $p\in S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. This completes the proof of the lemma. ∎
@@
$\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\geq\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}.$
@@
$\dim_{H}(S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\geq$ $\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
Using the same argument as in [17, Section 10], we conclude that
$\dim_{H}(S_{\tau}^{c})\geq$ $\dim\mathbf{G}(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
@@
Note that by Lemma 4.2, we have $\nu_{0}(a_{1})>0$ and the argument here works for all $0\leq\tau<\beta_{0}(a_{-1})$. This completes the proof of Theorem 1.3 (provided that $\{a_{t}\}_{t\in\mathbb{R}}$ is mixing on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$). ∎
5. Proof of Theorem 1.3: $\{a_{t}\}_{t\in\mathbb{R}}$ in a proper normal $\mathbb{Q}$-subgroup of $\mathbf{G}$
@@
In this section, we discuss the case when the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is not mixing, and explain how to modify the arguments in §3 and §4 and give a proof of Theorem 1.3.
@@
We first discuss the ergodic properties of group actions on homogeneous spaces, and one can refer to e.g. [29, 35, 37] for details. Let $\mathbf{G}_{i}$ $(1\leq i\leq k)$ be the $\mathbb{Q}$-simple factors of $\mathbf{G}$. Then $\mathbf{G}$ is an almost direct product of $\mathbf{G}_{i}$ $(1\leq i\leq k)$. Without loss of generality, we may assume that $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into $\mathbf{G}_{i}(\mathbb{R})^{0}$ ($1\leq i\leq s$) for some $s<k$. It is known that for each $1\leq i\leq k$, any arithmetic lattice $\Gamma_{i}$ inside $\mathbf{G}_{i}(\mathbb{Z})\cap\mathbf{G}_{i}(\mathbb{R})^{0}$ is irreducible in $\mathbf{G}_{i}(\mathbb{R})^{0}$ (and we fix such an arithmetic lattice $\Gamma_{i}$ for later use). Moreover, since $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into $\mathbf{G}_{i}(\mathbb{R})^{0}$ $(1\leq i\leq s)$, we have $\mathbf{T}\cap\mathbf{G}_{i}\neq\{e\}$ $(1\leq i\leq s)$. So $\mathbf{G}_{i}$ is $\mathbf{Q}$-isotropic, and $\mathbf{G}_{i}(\mathbb{R})^{0}/\Gamma_{i}$ is not compact. By Godement compactness criterion (Cf. [7, Theorem 11.6]) and [8, 6.21], every simple factor of $\mathbf{G}_{i}(\mathbb{R})^{0}$ is not compact. Let $\{a_{t}^{i}\}_{t\in\mathbb{R}}$ be the projection of $\{a_{t}\}_{t\in\mathbb{R}}$ in $\mathbf{G}_{i}(\mathbb{R})^{0}$ $(1\leq i\leq s)$. Then $\{a_{t}^{i}\}_{t\in\mathbb{R}}\subset\mathbf{T}(\mathbb{R})\cap\mathbf{G}_{i}(\mathbb{R})$ is a non-compact subgroup in $\mathbf{G}_{i}(\mathbb{R})^{0}$. Using Moore’s ergodicity theorem [33] and Mautner phenomenon [30, 34], one can conclude that the action of $\{a_{t}^{i}\}_{t\in\mathbb{R}}$ on $\mathbf{G}_{i}(\mathbb{R})^{0}/\Gamma_{i}$ is mixing $(1\leq i\leq s)$. Consequently, the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\prod_{i=1}^{s}\mathbf{G}_{i}(\mathbb{R})^{0}/\prod_{i=1}^{s}\Gamma_{i}$ is mixing.
@@
In the follwing, we fix these lattices $\Gamma_{i}$’s and $\Gamma$ and denote by
@@
Therefore, the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$ is mixing.
@@
which is a minimal parabolic $\mathbb{Q}$-subgroup of $\tilde{\mathbf{G}}$, and let $\overline{\tilde{\mathbf{P}}}_{0}=\overline{\mathbf{P}}_{0}\cap\tilde{\mathbf{G}}$. Let $U$ be an open bounded subset in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ and define
@@
In the following, we need results in the reduction theory about $\tilde{\mathbf{G}}(\mathbb{R})/\tilde{\Gamma}$. Let $\tilde{K}$ be a maximal compact subgroup in $\tilde{\mathbf{G}}(\mathbb{R})$. Let $\tilde{M}$ be the connected component of identity in the unique maximal $\mathbb{Q}$-anisotropic subgroup in $Z_{\tilde{\mathbf{G}}(\mathbb{R})}(\tilde{\mathbf{T}}(\mathbb{R}))$ (= the centralizer of $\tilde{\mathbf{T}}(\mathbb{R})$ in $\tilde{\mathbf{G}}(\mathbb{R})$). Denote by
@@
where $\tilde{\Delta}$ is the set of positive $\mathbb{Q}$-simple roots in $\tilde{\mathbf{G}}$. A Siegel set in $\tilde{\mathbf{G}}(\mathbb{R})$ is a subset of the form $\tilde{S}_{\eta,\Omega}=\tilde{K}\cdot\tilde{T}_{\eta}\cdot\tilde{\Omega}$ for some $\eta\in\mathbb{R}$ and a relatively compact open subset $\tilde{\Omega}$ containing identity in $\tilde{M}\cdot R_{u}(\tilde{\mathbf{P}}_{0})(\mathbb{R})$, and the group $\tilde{\mathbf{G}}(\mathbb{R})$ can be written as
@@
for some Siegel set $\tilde{S}_{\eta,\Omega}$ and some finite subset $\tilde{\mathcal{K}}\subset\tilde{\mathbf{G}}(\mathbb{Q})$. Moreover, the finite set $\tilde{\mathcal{K}}$ satisfies the property that
@@
Denote by $\tilde{\mathcal{K}}=\{\tilde{x}_{j}\}_{j\in\tilde{J}}\subset\tilde{\mathbf{G}}(\mathbb{Q})$ and we may assume that $e\in\tilde{\mathcal{K}}$.
@@
    $g\in R_{u}(\overline{\tilde{\mathbf{P}}}_{0})\cap(\tilde{\mathbf{H}}_{1}(\mathbb{R})\cdot\tilde{\mathbf{P}}_{0}(\mathbb{R})\cdot\tilde{\mathcal{K}}\cdot\tilde{\Gamma}).$
@@
Let $\tilde{\mathfrak{a}}$ be the Lie algebra of $\tilde{\mathbf{T}}(\mathbb{R})$. For any $\tilde{x}_{j}\in\tilde{\mathcal{K}}$, any compact subset $K_{1}$ in $\ker(\beta_{0})\cap\tilde{\mathfrak{a}}$, any compact subset $K_{2}$ in $\tilde{\mathbf{H}}_{e}(\mathbb{R})$, any compact subset $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}$ and any compact subset $K_{4}\subset R_{u}(\tilde{\mathbf{P}}_{0})(\mathbb{R})$, we define
@@
to be the set of all rational points $q$ in $U$ such that
@@
    $q=a\cdot h\cdot m\cdot u\cdot\tilde{x}_{j}\cdot\gamma$ for some $a\in\exp(\tilde{\mathfrak{a}})$, $\pi_{\ker(\beta_{0})}(a)\in K_{1}$, $h\in K_{2}$, $m\in K_{3}$, $u\in K_{4}$ and $\gamma\in\tilde{\Gamma}$.
@@
$\tilde{\mathbf{H}}_{e}\times\{q\}\to\tilde{\mathbf{H}}_{e}\cdot q\subset R_{u}(\overline{\tilde{\mathbf{P}}}_{0}).$
@@
Let $U$ be a small open bounded subset in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})$. Then for any $j\in\tilde{J}$ with $\tilde{x}_{j}\in\tilde{\mathbf{G}}(\mathbb{R})^{0}$ and any sufficiently large $l>0$, we have
$m_{\tilde{\mathbf{H}}_{e}}(\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\tilde{\mathbf{P}}}_{0})}(U)$
@@
$m_{\tilde{\mathbf{H}}_{e}}(\tilde{S}_{K_{1},K_{2},K_{3},K_{4}}(U,A_{l},B_{l}))\sim l^{-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\tilde{\mathbf{P}}}_{0})}(U).$
@@
The analogues of Lemma 4.1 and Lemma 4.2 hold as well in $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$.
@@
for some $T>0$ and $l>1$. Let ~$\tilde{\mathcal{F}}_{q}=\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot q$ be the leaf through $q\in\tilde{\mathbf{F}}_{e}(\mathbb{Q})$ such that
~$\tilde{\mathcal{F}}_{q}\cap\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\neq\emptyset.$

### v2

@@ 3. Counting rational elements
$a_{-t_{k}}\cdot B_{\mathbf{F}_{e}}(\tilde{\delta}_{0})\cdot a_{t_{k}}\subset B_{\mathbf{F}_{e}}(\tilde{\delta}_{0},T_{k})\textup{ and }a_{T_{k}}\cdot q\Gamma\in\exp(\mathfrak{a}_{I_{0},K_{1}})\tilde{K}_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma\;(\forall q\in\mathcal{H}_{k,S}).$
@@
if $T_{k}>0$ and $l_{k}>0$ are sufficiently large. Note that $\Phi^{0}(\mathbf{F}_{e})$ may not be empty, and some of the cubes in $\mathcal{P}_{k,S}$ may be outside $S$, and here we collect only those cubes in $\mathcal{P}_{k,S}$ which are inside $S$. Thus, we obtain a family of disjoint cubes inside $S$ constructed from $\mathcal{P}_{k,S}$. We denote by $\mathcal{A}_{k}$ the collection of all the disjoint cubes inside $S$ constructed from $\mathcal{P}_{k,S}$ where $S$ ranges over all elements in $\mathcal{A}_{k-1}$.
In this manner, we obtain an increasing sequence of sufficiently large numbers $\left\{l_{k}\right\}_{k\in\mathbb{N}}$ and a tree-like collection $\mathcal{A}=\left\{\mathcal{A}_{k}\right\}_{k\in\mathbb{N}}$ of cubes in $U_{0}$. Using the notation in Theorem 4.2, we have
$d_{k}(\mathcal{A})\asymp l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}$
@@
$\begin{cases}\Delta_{k}(\mathcal{A})\asymp l_{k+1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\beta_{0}(a_{1})}}\cdot\prod_{\alpha\in\Phi(\mathbf{F}_{e})}l_{k+1}^{\frac{\alpha(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}},&\Phi^{0}(\mathbf{F}_{e})=\emptyset\\ \Delta_{k}(\mathcal{A})\asymp l_{k+1}^{-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\beta_{0}(a_{1})}}\cdot\prod_{\alpha\in\Phi^{1}(\mathbf{F}_{e})}l_{k+1}^{\frac{\alpha(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}}\cdot\prod_{\alpha\in\Phi^{0}(\mathbf{F}_{e})}l_{k}^{\frac{\nu_{0}(a_{-1})}{\beta_{0}(a_{-1})-(\tau+\epsilon)}},&\Phi^{0}(\mathbf{F}_{e})\neq\emptyset.\end{cases}$
Now let $\mathbf{A}_{\infty}=\bigcap_{k\in\mathbb{N}}\mathbf{A}_{k}$. By Theorem 4.2, we can compute (assuming that $l_{k+1}$ is much larger than $l_{k}$ for any $k\in\mathbb{N}$) that
@@
$=$ $\dim R_{u}(\overline{\mathbf{P}}_{0})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})(\tau+\epsilon)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}.$
By the definition of $\tau=\tau(\psi)$, there exists a divergent sequence $\{s_{k}\}_{k\in\mathbb{N}}\subset\mathbb{R}_{+}$ such that
$\tau=\tau(\psi)=\lim_{k\to\infty}\left(-\frac{\ln\psi(s_{k})}{s_{k}}\right).$
Now we choose the sequence $\{l_{k}\}_{k\in\mathbb{N}}$ in the construction of $\mathbf{A}_{\infty}$ so that the sequence $\{t_{k}\}_{k\in\mathbb{N}}$ equals the sequence $\{s_{k}\}_{k\in\mathbb{N}}$ defined as above. Then we can prove the following
@@
We have $\mathbf{A}_{\infty}\subset S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$.
@@
By the construction, for any $p\in\mathbf{A}_{\infty}$, there exists a sequence of rational elements
@@
Note that by definition, the height of the rational element $a_{t_{k}}\cdot q_{k}$ is equal to the co-volume of the lattice $\rho(a_{t_{k}}\cdot q_{k})\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}$ and
$0pt(a_{t_{k}}\cdot q_{k})=e^{\beta_{0}(a_{t_{k}})\cdot\dim V_{\beta_{0}}}\cdot d(q_{k}).$
By Lemma 2.6, we deduce that for any $k\in\mathbb{N}$
$\delta(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})$ $\asymp\delta(\rho(a_{t_{k}}\cdot q_{k})\mathbb{Z}^{d})\lesssim 0pt(a_{t_{k}}\cdot q_{k})^{\frac{1}{\dim V_{\beta_{0}}}}$
$\lesssim e^{\beta_{0}(a_{t_{k}})}\cdot l_{k}\leq e^{-(\tau+\epsilon)t_{k}}.$
Note that by our choice of $\{t_{k}\}_{k\in\mathbb{N}}$, for $\epsilon>0$ and for sufficiently large $k\in\mathbb{N}$,
$\frac{\ln\psi(t_{k})}{t_{k}}\leq\tau+\frac{\epsilon}{2}\implies\psi(t_{k})\geq e^{-(\tau+\epsilon/2)t_{k}}.$
So we have
$\delta(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})$ $\lesssim e^{-(\tau+\epsilon)t_{k}}\leq\psi(t_{k})\cdot e^{-\frac{\epsilon}{2}t_{k}}$
and by definition $p\in S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. This completes the proof of the lemma. ∎
@@
$\dim_{H}(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\geq\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})(\tau+\epsilon)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}.$
@@
$\dim_{H}(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\geq$ $\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}.$
By definition, if an element $g\in\mathbf{G}(\mathbb{R})$ belongs to $S_{\rho}(\psi)^{c}$, then for any $h\in\mathbf{P}_{0}(\mathbb{R})$, $h\cdot g$ also belongs to $S_{\rho}(\psi)^{c}$. Using the same argument as in [25, §10], we can conclude that
$\dim_{H}(S_{\rho}(\psi)^{c})\geq$ $\dim\mathbf{P}_{0}(\mathbb{R})+\dim_{H}(S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\geq$ $\dim\mathbf{G}(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}.$
@@
This completes the proof of Theorem 1.4 in the case where $0\leq\tau(\psi)<\beta_{0}(a_{-1})$ and $\{a_{t}\}_{t\in\mathbb{R}}$ is mixing on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. ∎
5. Proof of Theorem 1.4: The other cases
@@
In this section, we discuss the case where $0\leq\tau(\psi)<\beta_{0}(a_{-1})$ and the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is not mixing, and explain how to modify the arguments in §3 and §4 to prove Theorem 1.4. We study the case $\tau(\psi)=\beta_{0}(a_{-1})$ at the end of this section.
@@
We first discuss the ergodic properties of group actions on homogeneous spaces, for which one may refer to [59, 66, 69]. Let $\mathbf{G}_{i}$ $(1\leq i\leq k)$ be the $\mathbb{Q}$-simple factors of $\mathbf{G}$. Then $\mathbf{G}$ is an almost direct product of $\mathbf{G}_{i}$ $(1\leq i\leq k)$. Without loss of generality, we may assume that $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into $\mathbf{G}_{i}(\mathbb{R})^{0}$ ($1\leq i\leq s$) for some $s<k$. It is known that for each $1\leq i\leq k$, any arithmetic lattice $\Gamma_{i}$ inside $\mathbf{G}_{i}(\mathbb{Z})\cap\mathbf{G}_{i}(\mathbb{R})^{0}$ is irreducible in $\mathbf{G}_{i}(\mathbb{R})^{0}$ (and we fix such an arithmetic lattice $\Gamma_{i}$ for later use). Moreover, since $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into $\mathbf{G}_{i}(\mathbb{R})^{0}$ $(1\leq i\leq s)$, we have $\mathbf{T}\cap\mathbf{G}_{i}\neq\{e\}$ $(1\leq i\leq s)$. So $\mathbf{G}_{i}$ is $\mathbf{Q}$-isotropic, and $\mathbf{G}_{i}(\mathbb{R})^{0}/\Gamma_{i}$ is not compact. By Godement compactness criterion (Cf. [9, Theorem 11.6]) and [10, 6.21], every simple factor of $\mathbf{G}_{i}(\mathbb{R})^{0}$ is not compact. Let $\{a_{t}^{i}\}_{t\in\mathbb{R}}$ be the projection of $\{a_{t}\}_{t\in\mathbb{R}}$ in $\mathbf{G}_{i}(\mathbb{R})^{0}$ $(1\leq i\leq s)$. Then $\{a_{t}^{i}\}_{t\in\mathbb{R}}\subset\mathbf{T}(\mathbb{R})\cap\mathbf{G}_{i}(\mathbb{R})$ is a non-compact subgroup in $\mathbf{G}_{i}(\mathbb{R})^{0}$. Using Moore’s ergodicity theorem [64] and Mautner phenomenon [60, 65], one can conclude that the action of $\{a_{t}^{i}\}_{t\in\mathbb{R}}$ on $\mathbf{G}_{i}(\mathbb{R})^{0}/\Gamma_{i}$ is mixing $(1\leq i\leq s)$. Consequently, the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\prod_{i=1}^{s}\mathbf{G}_{i}(\mathbb{R})^{0}/\prod_{i=1}^{s}\Gamma_{i}$ is mixing.
@@
In the follwing, we fix the lattices $\Gamma_{i}$’s and $\Gamma$ and denote by
@@
So the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$ is mixing.
@@
which is a minimal parabolic $\mathbb{Q}$-subgroup of $\tilde{\mathbf{G}}$, and let $\overline{\tilde{\mathbf{P}}}_{0}=\overline{\mathbf{P}}_{0}\cap\tilde{\mathbf{G}}$. Let $U$ be a small open bounded subset in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R})$ and define
@@
[…]
@@
[…]
@@
[…]
@@
[…]
The proof of the following lemma is similar to those of Corollary 2.3 and Lemma 2.4.
@@
[…]
@@
[…]
@@
to be the set of all rational elements $q$ in $U$ such that
@@
[…]
@@
$\tilde{\mathbf{H}}_{e}(\mathbb{R})\times\{q\}\to\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot q\subset R_{u}(\overline{\tilde{\mathbf{P}}}_{0})(\mathbb{R}).$
@@
[…]
$m_{\tilde{\mathbf{H}}_{e}}(\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\asymp l^{-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\tilde{\mathbf{P}}}_{0})}(U)$
@@
$m_{\tilde{\mathbf{H}}_{e}}(\tilde{S}_{K_{1},K_{2},K_{3},K_{4}}(U,A_{l},B_{l}))\asymp l^{-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\alpha(a_{1})/\beta_{0}(a_{1})}\cdot\mu_{R_{u}(\overline{\tilde{\mathbf{P}}}_{0})}(U).$
@@
The analogue of Lemma 4.1 holds as well in $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$.
@@
for some $T>0$ and $l>1$. Let $\tilde{\mathcal{F}}_{q}=\tilde{\mathbf{H}}_{e}(\mathbb{R})\cdot q$ be the leaf through $q\in\tilde{\mathbf{F}}_{e}(\mathbb{Q})$ such that
$\tilde{\mathcal{F}}_{q}\cap\tilde{S}_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\neq\emptyset.$